(a) Suppose f and g are functions whose domains are subsets of z+, the set of positive integers. Give the definition of 'f is "\Omicron"(g)'


(b) Use the definition of 'f is "\Omicron"(g)' to show that:

(I) 2n+27 is "\Omicron"(3n)

(ii) 5n is not "\Omicron"(4n)


1
Expert's answer
2021-07-28T09:58:57-0400

Let "f" and "g" are functions whose domains are subsets of "\Z_+", the set of positive integers. We say that "f" is "O(g)" if there exist a positive real number "c" and a positive integer "n_0" such that "f(n)\le cg(n)" for all "n\ge n_0."


(b) Let us use the definition of 'f is "\Omicron"(g)' to show that the following statements.


(I) Since "2^n+27<3^n+27\cdot 3^n=28\cdot 3^n" for each positive integer "n," we put "c=28" and conclude that "2^n+27" is "O(3^n)."


(ii) Let us show that "5^n" is not "O(4^n)." Let us prove using the method by contradiction. Suppose that there exist a positive real number "c" and a positive integer "n_0" such that "5^n\le c\cdot 4^n" for all "n\ge n_0."

It follows that "(\frac{5}{4})^n\le c" for all "n\ge n_0." Since "\lim\limits_{n\to+\infty}(\frac{5}{4})^n=+\infty," we conclude that there exist "m\in\Z_+" such that "(\frac{5}{4})^n>c" for all "n\ge m." This contradiction proves the statement.